OpenStudy (anonymous):

A spherical ballon is being filled with helium at a rate of 8cm^3/s. At what rate is its radius increasing, when it has been filling for 50 seconds?

OpenStudy (bahrom7893):

dV/dt = 8cm^3/s t = 50s dr/dt - ?

OpenStudy (bahrom7893):

V = 4/3*pie*r^3 dV/dt = 4*pie*r^2

OpenStudy (bahrom7893):

woops sorry forgot dr/dt

OpenStudy (bahrom7893):

dV/dt = 4*pie*r^2*(dr/dt)

OpenStudy (bahrom7893):

8 = 4*pie*r^2*(dr/dt) 2/pie = r^2*(dr/dt)

OpenStudy (bahrom7893):

was that all info given?

OpenStudy (bahrom7893):

was it??